# spacing=nonfrench

### v1

@@ 1.3 Plemelj-Calderón Problem
This defines a function which is holomorphic outside the curve $\Gamma$. Moreover, this function has boundary values on $\Gamma$ from inside and outside, and one has Plemelj formula on $\Gamma$:
@@
where the Cauchy integral
@@
Of course the result remains true for much more general curves and functions. Subsequent generalizations relax the smoothness requirements on the curve $\Gamma$ and the function $f$: In order for Plemelj formula to hold true one indeed only needs $\Gamma$ to be rectifiable and $f\in L^{1}(\Gamma)$. The modern approach of this problem is then to solve the problem in a given space: Calderón ([6]) asked for instance for which rectifiable curves it is true that for any $f\in L^{2}(\Gamma)$ Plemelj functions $F_{i}|_{\Gamma}$ and $F_{e}|_{\Gamma}$ are also in $L^{2}(\Gamma)$? Equivalently, when is the operator $T$ bounded on $f\in L^{2}(\Gamma)$? This problem has been solved by Coifman-McIntosh-Meyer ([10]) for Lipschitz curves (Calderón had previously solved it for curves with small Lipschitz bound ([6])) and then the last word was given by David ([11]) who proved that $T$ is bounded on $f\in L^{2}(\Gamma)$ (or $L^{p}(\Gamma)$, $1<p<\infty$) if and only if $\Gamma$ is Ahlfors-regular, meaning that the length of the part of $\Gamma$ lying in a disk of radius $r$ is no bigger than $Cr$ for some $C$ independent of the chosen disk.
Now Plemelj-Calderón problem may be addressed for curves that are not necessarily rectifiable. An example has been given in [48] where Plemelj-Calderón problem for Hölder classes has been raised and solved in some particular cases. The idea is to replace the Cauchy integral by the identity
@@
which is true for any test function $f\in C_{c}^{\infty}(\mathbb{C})$, and from which it follows that, in the case of rectifiable boundary,
$F_{i}(z)$ $=-\frac{1}{\pi}\iint_{\Omega_{e}}\frac{\bar{\partial}f(\zeta)}{\zeta-z}d\xi d\eta,\qquad z\in\Omega_{i};$
$F_{e}(z)$ $=\frac{1}{\pi}\iint_{\Omega_{i}}\frac{\bar{\partial}f(\zeta)}{\zeta-z}d\xi d\eta,\qquad z\in\Omega_{e}$
where $\Omega_{i,e}$ stand for the interior and exterior components of $\Gamma$. Later Astala ([3]) obtained more precise results for the Hölder classes: these will be discussed in section 3 where we will draw the parallel with fractional Sobolev spaces.
In this paper we address the following Plemelj-Calderón problem: given $p>1,\,s\in(0,1)$, can one write a function $f\in\mathcal{B}_{p,p}^{s}(\Gamma)$ as a sum of two functions $F_{i,e}$ that are holomorphic in $\Omega_{i,e}$ and whose boundary values belong to $\mathcal{B}_{p,p}^{s}(\Gamma)$?
The main result of the present paper in this direction is Theorems 2.3, 2.4, 3.5 and Corollaries 3.8, 3.17. A priori version of these results is as follows.
@@
Let $\Gamma$ be a quasicircle , $p>1,\,s\in(0,1)$. Then a priori Plemelj-Calderón property holds for $C_{c}^{\infty}(\mathbb{C})|_{\Gamma}$, if
$(p-1)(h(\Gamma)-1)<sp<p+1-h(\Gamma).$
2 Critical Besov Spaces on Quasicircles
Recall that a quasidisk is the image of the unit disk $\mathbb{D}_{i}$ by a quasiconformal homeomorphism of the plane, and its boundary curve is called a quasicircle. An homeomorphism $\Phi$ of the plane is called quasiconformal if its gradient in the sense of distribution is in $L^{2}_{loc}(\mathbb{C})$ and if in addition there exists a constant $k<1$ such that
@@ 2 Critical Besov Spaces on Quasicircles
Any quasicircle $\Gamma$ admits a quasiconformal reflection across $\Gamma$ which is bi-Lipschitz around $\Gamma$. Here, by a quasiconformal reflection we mean an anti-quasiconformal self-homeomorphism $R$ of $\overline{\mathbb{C}}$ such that $R\circ R=\text{id}$ on $\overline{\mathbb{C}}$ and $R=\text{id}$ when restricted to $\Gamma$. Precisely, let $\Phi$ be a quasiconformal homeomorphism of $\overline{\mathbb{C}}$ we then have $R(z)=\Phi\circ(1/\overline{\Phi^{-1}(z)})$ is a quasiconformal reflection with respect to the quasicircle $\Gamma:=\partial\Phi(\mathbb{D}_{i})$.
Let $\Omega_{i}$ and $\Omega_{e}$ be domains bounded by a quasicircle $\Gamma$. If $\varphi_{i}$ and $\varphi_{e}$ are conformal isomorphisms from $\mathbb{D}_{i}$ onto $\Omega_{i}$ and from $\mathbb{D}_{e}$ onto $\Omega_{e}$, respectively, then they can be extended to a homeomorphism between the closures such that $h:=\varphi_{e}^{-1}\circ\varphi_{i}$ is a homeomorphism of $\mathbb{T}$ which is quasisymmetric in the sense that there exists a constant $C>0$ such that
@@
Let $\Gamma$ be a Jordan curve and $p>1$. We recall that the spaces $\mathcal{B}_{p,p}^{1/p}(\Omega_{i,e})$ , called the $p$-critical Besov spaces, have the remarkable property of being conformally invariant: if $\Omega,\Omega^{\prime}$ are two Jordan domains and $\varphi:\Omega\to\Omega^{\prime}$ is a biholomorphism then the map $f\mapsto f\circ\varphi$ is a quasi-isometry between $\mathcal{B}_{p,p}^{1/p}(\Omega^{\prime})$ and $\mathcal{B}_{p,p}^{1/p}(\Omega)$. In order to be as much as possible self-contained, here is a sketch of proof: by the change of variable $\zeta=\varphi(z)$ we have
@@
Now the Koebe distortion theorem [1] implies that
@@
Applying this result for $\Omega^{\prime}=\mathbb{D}$, $\partial\mathbb{D}$ being the unit circle $\mathbb{T}$, we see that all the $p$-critical Besov spaces over $\Omega$ are quasi-isometric to $\mathcal{B}_{p,p}^{1/p}(\mathbb{D})$. It is a classical result [37] that the space $\mathcal{C}_{p}(\mathbb{D})$ consisting of functions in $\mathcal{B}_{p,p}^{1/p}(\mathbb{D})$ that have a continuous extension to $\overline{\mathbb{D}}$ is dense in $\mathcal{B}_{p,p}^{1/p}(\mathbb{D})$. This can also be derived from that the set of polynomials is dense in the Bergman space $A^{p}_{p-2}$, the set of analytic functions $g$ in $\mathbb{D}$ such that
@@
Since a bi-holomorphic homeomorphism between two Jordan domains extends to a homeomorphism between the closures the same property will hold for the analogous spaces $\mathcal{C}_{p}(\Omega)$.
The next step is to prove that $\mathcal{C}_{p}(\Omega_{i})|_{\Gamma}=\mathcal{C}_{p}(\Omega_{e})|_{\Gamma}$ for any quasicircle $\Gamma$. We first prove the result for $p=2$.
Let $u_{i}\in\mathcal{C}_{2}(\Omega_{i})$ and $f=u_{i}|_{\Gamma}$. Recall that there exists a quasiconformal reflection $R$ across $\Gamma$. Put $v(z)=u_{i}(R(z)),z\in\Omega_{e}$. By the bi-Lipschitz property of $R$ around $\Gamma$ we have that
@@
(stricly speaking, there is a technical problem at $\infty$, which is settled in the appendix). We invoke then the Dirichlet principle which implies that
@@
where $u_{e}:=P_{e}(f)$ stands for the harmonic extension of $f$ in $\Omega_{e}$. That is, we have that the transmission operator
@@
If $p>1$ and $0<s<1$ (except the case of $p=2$ and $s=1/2$) we loose the Dirichlet principle and use instead the almost-Dirichlet principle. Recall that the almost-Dirichlet principle holds for $\mathcal{B}_{p,p}^{s}(\Omega)$ if for any $f$ defined on $\Gamma$, $u$ denotes its harmonic extension in $\Omega$ and $v$ is any continuous extension of $f$ to $\Omega$ which is in $C^{1}(\Omega)$, we have
@@
for some constant $C$ depending only on $\Omega,\,p,\,s$. It was proved in [30] that the almost-Dirichlet principle holds for $\mathcal{B}_{p,p}^{s}(\mathbb{D}_{i,e})$ for all p>1, 0<s<1p>1,\,0<s<1.
@@
As a consequence, if $f$ is a continuous function on $\Gamma$, its harmonic extension to $\Omega_{i}$ is in $\mathcal{C}_{p}(\Omega_{i})$ if and only if its harmonic extension to $\Omega_{e}$ is in $\mathcal{C}_{p}(\Omega_{e})$. Then, $C_{p}(\Omega_{i}\!\to\!\Gamma)=C_{p}(\Omega_{e}\!\to\!\Gamma)$ so that we may define $\mathcal{C}_{p}(\Gamma):=\mathcal{C}_{p}(\Omega_{i,e}\!\to\!\Gamma)$.
@@
Let $\varphi_{i,e}$ denote the Riemann mappings from $\mathbb{D}_{i,e}$ onto $\Omega_{i,e}$ as before. By the conformal invariance, we have that $f\circ\varphi_{i,e}\in\mathcal{C}_{p}(\mathbb{T})$ and $f\circ\varphi_{e}=f\circ\varphi_{i}\circ h$ with $h:=\varphi_{i}^{-1}\circ\varphi_{e}$ being a circle homeomorphism called conformal welding. As we have recalled above this homeomorphism is a quasisymmetry: we deduce from this that $V_{h}:g\mapsto g\circ h$ is an isomorphism of $\mathcal{C}_{p}(\mathbb{T})$ which extends to an isomorphism of $\mathcal{B}_{p,p}^{1/p}(\mathbb{T})$ with $V_{h}^{-1}=V_{h^{-1}}$. Now if one consider any quasisymmetry $h$ of the unit circle we know [1] that $h$ is the conformal welding of some quasicircle. We deduce from this geometric approach that $V_{h}$ is an isomorphism of $\mathcal{B}_{p,p}^{1/p}(\mathbb{T})$ for any quasisymmetry $h$. Conversely, Nag-Sullivan [33] for the case $p=2$ and Bourdaud [5] for the general case have proven that if $h$ is a circle homeomorphism and if $V_{h}$ is an isomorphism of $\mathcal{B}_{p,p}^{1/p}(\mathbb{T})$ then $h$ must be a quasisymmetry. Vodop’Yanov[42], and later Bourdaud [5] moreover proved that if $s\neq 1/p$ then the circle homeomorphisms $h$ such that $V_{h}$ operates on $\mathcal{B}_{p,p}^{s}(\mathbb{T})$ isomorphically are the bi-Lipschitz ones which form a proper subgroup of the group of quasisymmetries.
2.1 Boundary Values of Functions $\mathcal{B}_{p,p}^{1/p}(\Omega)$

### v2

@@ 1.3 Plemelj-Calderón problem: statement of two more main theorems
This defines a function that is holomorphic outside the curve $\Gamma$. Moreover, this function has boundary values on $\Gamma$ from inside and outside, and one has Plemelj formulae on $\Gamma$:
@@
where $F_{i}=F|_{\Omega_{i}}$, $F_{e}=F|_{\Omega_{e}}$ and the Cauchy integral
@@
Of course, the result remains true for much more general curves and functions. Subsequent generalizations relax the smoothness requirements on the curve $\Gamma$ and the function $f$: In order for Plemelj formulae to hold, one indeed only needs $\Gamma$ to be rectifiable and $f\in L^{1}(\Gamma)$. The modern approach of this problem is then to solve the problem in a given space: Calderón ([6]) asked, for instance, for which rectifiable curves it is true that for any $f\in L^{2}(\Gamma)$ Plemelj functions $F_{i}|_{\Gamma}$ and $F_{e}|_{\Gamma}$ are also in $L^{2}(\Gamma)$? Equivalently, when is the operator $T$ bounded on $f\in L^{2}(\Gamma)$? This problem has been solved by Coifman-McIntosh-Meyer ([10]) for Lipschitz curves (Calderón had previously solved it for curves with small Lipschitz bound ([6])) and then the last word was given by David ([11]) (using the Lipschitz result [10]) who proved that $T$ is bounded on $f\in L^{2}(\Gamma)$ (or $L^{p}(\Gamma)$, $1<p<\infty$) if and only if $\Gamma$ is Ahlfors-regular, meaning that the length of the part of $\Gamma$ lying in a disk of radius $r$ is no bigger than $Cr$ for some $C$ independent of the chosen disk.
Now Plemelj-Calderón problem may be addressed for curves that are not necessarily rectifiable. An example has been given in [54] where Plemelj-Calderón problem for Hölder classes has been raised and solved in some particular cases. The idea is to replace the Cauchy integral with the identity
@@
which is true for any test function $f\in C_{c}^{\infty}(\mathbb{C})$, and from which it follows that, in the case of a rectifiable boundary,
$F_{i}(z)$ $=-\frac{1}{\pi}\iint_{\Omega_{e}}\frac{\bar{\partial}f(\zeta)}{\zeta-z}d\xi d\eta,\qquad z\in\Omega_{i},$
$F_{e}(z)$ $=\frac{1}{\pi}\iint_{\Omega_{i}}\frac{\bar{\partial}f(\zeta)}{\zeta-z}d\xi d\eta,\qquad z\in\Omega_{e}.$
Later, Astala ([3]) obtained more precise results for the Hölder classes: these will be discussed in Section 2 where we will draw a parallel with fractional Besov-Sobolev spaces $\mathcal{B}_{p,p}^{s}(\Gamma)$. Recently, Plemelj-Calderón problem for critical Besov spaces (the case of $s=1/p$) has been treated by Liu-Shen [28], and by Matsuzaki [31]. Schippers-Staubach has further discussed this issue in a broader sense; see [42] and references therein.
In Section 2, we address the following Plemelj-Calderón problem: given $p>1,\,s\in(0,1)$, can one write a function $f\in\mathcal{B}_{p,p}^{s}(\Gamma)$ as a sum of two functions $F_{i,e}$ that are holomorphic in $\Omega_{i,e}$ and whose boundary values belong to $\mathcal{B}_{p,p}^{s}(\Gamma)$?
The main result of the present paper in this direction is Theorems 2.3, 2.4, 2.9 and Corollaries 2.13, 3.6 (each focusing on different $(p,s)$-values). The a priori version of these results is as follows.
@@
Suppose that $\Gamma$ is a quasicircle. For $p>1,\,s\in(0,1)$, the a priori Plemelj-Calderón property holds for $C_{c}^{\infty}(\mathbb{C})|_{\Gamma}\cap\mathcal{B}_{p,p}^{s}(\Gamma)$ if
$(p-1)(h(\Gamma)-1)<sp<p+1-h(\Gamma)$
that reduces to $0<s<1$ when $\Gamma$ is chord-arc.
Assume that the curve $\Gamma$ is rectifiable so that the space $B_{p,p}^{s}(\Gamma)$ is well-defined. We will see that Plemelj-Calderón problem on $B_{p,p}^{s}(\Gamma)$ is much easier to handle: In Section 4, we will show the following:
Theorem 1.6 (see Theorem 4.4).
Suppose that the curve $\Gamma$ is chord-arc. For $0<s<1$ and $1<p<\infty$, the Cauchy integral operator $T$ is bounded on $B_{p,p}^{s}(\Gamma)$.
2 Plemelj-Calderón problem for $\mathcal{B}_{p,p}^{s}(\Gamma)$ on quasicircles: the non-rectifiable case
The main goal of this section is to address Plemelj-Calderón Problem for the space $\mathcal{B}_{p,p}^{s}(\Gamma)$, $p>1$ and $0<s<1$, on quasicircles $\Gamma$. We now introduce several notions closely related to quasicircles which will be used later.
Recall that a quasidisk is the image of the unit disk $\mathbb{D}_{i}$ under a quasiconformal homeomorphism of the plane, and its boundary curve is called a quasicircle. An homeomorphism $\Phi$ of the plane is called quasiconformal if its gradient in the sense of distribution is in $L^{2}_{loc}(\mathbb{C})$ and if, in addition, there exists a constant $k<1$ such that
@@
Any quasicircle $\Gamma$ admits a quasiconformal reflection across $\Gamma$ that is bi-Lipschitz around $\Gamma$. Here, by a quasiconformal reflection we mean an anti-quasiconformal self-homeomorphism $R$ of $\overline{\mathbb{C}}$ such that $R\circ R=\text{id}$ on $\overline{\mathbb{C}}$ and $R=\text{id}$ when restricted to $\Gamma$. Precisely, let $\Phi$ be a quasiconformal homeomorphism of $\overline{\mathbb{C}}$ we then have $R(z)=\Phi\circ(1/\overline{\Phi^{-1}(z)})$ is a quasiconformal reflection with respect to the quasicircle $\Gamma=\partial\Phi(\mathbb{D}_{i})$.
Let $\Omega_{i}$ and $\Omega_{e}$ be domains bounded by a quasicircle $\Gamma$. If $\varphi_{i}$ and $\varphi_{e}$ are conformal isomorphisms from $\mathbb{D}_{i}$ onto $\Omega_{i}$ and from $\mathbb{D}_{e}$ onto $\Omega_{e}$, respectively, then they can be extended to a homeomorphism between the closures such that $h=\varphi_{e}^{-1}\circ\varphi_{i}$ is a homeomorphism of $\mathbb{T}$ that is quasisymmetric in the sense that there exists a constant $C>0$ such that
@@
2.1 Boundary values of functions in a dense subspace of $\mathcal{B}_{p,p}^{1/p}(\Omega)$
Let $\Gamma$ be a Jordan curve and $p>1$. We recall that the spaces $\mathcal{B}_{p,p}^{1/p}(\Omega_{i,e})$ , called the $p$-critical Besov spaces, have the remarkable property of being conformally invariant: if $\Omega,\Omega^{\prime}$ are two Jordan domains and $\varphi:\Omega\to\Omega^{\prime}$ is a biholomorphism, then the map $u\mapsto u\circ\varphi$ is a quasi-isometry between $\mathcal{B}_{p,p}^{1/p}(\Omega^{\prime})$ and $\mathcal{B}_{p,p}^{1/p}(\Omega)$. In order to be as self-contained as possible, here is a sketch of the proof: by the change of variables $\zeta=\varphi(z)$ we have
@@ 2.1 Boundary values of functions in a dense subspace of $\mathcal{B}_{p,p}^{1/p}(\Omega)$
Now, the Koebe distortion theorem (see [1]) implies that
@@
Applying this result to $\Omega^{\prime}=\mathbb{D}$, the inner or outer disk bounded by the unit circle $\mathbb{T}$, we see that all the $p$-critical Besov spaces over $\Omega$ are quasi-isometric to $\mathcal{B}_{p,p}^{1/p}(\mathbb{D})$. It is a classical result [44] that the space $\mathcal{C}_{p}(\mathbb{D})$ consisting of functions in $\mathcal{B}_{p,p}^{1/p}(\mathbb{D})$ that have a continuous extension to $\overline{\mathbb{D}}$ is dense in $\mathcal{B}_{p,p}^{1/p}(\mathbb{D})$. This can also be derived from the fact that the set of polynomials is dense in the Bergman space $A^{p}_{p-2}$, the set of analytic functions $g$ in $\mathbb{D}$ such that
@@
Since a bi-holomorphic homeomorphism between two Jordan domains extends to a homeomorphism between the closures, the same property will hold for the analogous space $\mathcal{C}_{p}(\Omega)$.
In this subsection, we prove that $\mathcal{C}_{p}(\Omega_{i})|_{\Gamma}=\mathcal{C}_{p}(\Omega_{e})|_{\Gamma}$ for any quasicircle $\Gamma$. To proceed, we first address the case of $p=2$.
Let $u_{i}\in\mathcal{C}_{2}(\Omega_{i})$ and $f=u_{i}|_{\Gamma}$. Recall that there exists a quasiconformal reflection $R$ across $\Gamma$. Put $v(z)=u_{i}(R(z)),z\in\Omega_{e}$. By the bi-Lipschitz property of $R$ around $\Gamma$ we have the following
@@
(strictly speaking, there is a technical problem at $\infty$, which is settled in the appendix). We then invoke the Dirichlet principle which implies that
@@
where $u_{e}=P_{e}(f)$ stands for the harmonic extension of $f$ in $\Omega_{e}$. That is, we have that the transmission operator
@@
For the case of a general $p\neq 2$ we loose the Dirichlet principle and use instead the almost-Dirichlet principle. Recall that the almost-Dirichlet principle is said to be valid for $\mathcal{B}_{p,p}^{s}(\Omega)$, p>1, 0<s<1p>1,\,0<s<1, if for any $f$ defined on $\Gamma$, $u$ denotes its harmonic extension in $\Omega$ and $v$ is any continuous extension of $f$ to $\Omega$ which is in $C^{1}(\Omega)$, we have
@@
for some constant $C$ depending only on $\Omega,\,p,\,s$. In [33] it was proved that the almost-Dirichlet principle holds for $\mathcal{B}_{p,p}^{s}(\mathbb{D}_{i,e})$ for all p>1, 0<s<1p>1,\,0<s<1.
@@
As a consequence, if $f$ is a continuous function on $\Gamma$, its harmonic extension to $\Omega_{i}$ is in $\mathcal{C}_{p}(\Omega_{i})$ if and only if its harmonic extension to $\Omega_{e}$ is in $\mathcal{C}_{p}(\Omega_{e})$. Then, $C_{p}(\Omega_{i}\!\to\!\Gamma)=C_{p}(\Omega_{e}\!\to\!\Gamma)$ so that we may define $\mathcal{C}_{p}(\Gamma)=\mathcal{C}_{p}(\Omega_{i,e}\!\to\!\Gamma)$.
@@
Let $\varphi_{i,e}$ denote the Riemann mappings from $\mathbb{D}_{i,e}$ onto $\Omega_{i,e}$ as before. By conformal invariance, we have $f\circ\varphi_{i,e}\in\mathcal{C}_{p}(\mathbb{T})$ and $f\circ\varphi_{e}=f\circ\varphi_{i}\circ h$ with $h=\varphi_{i}^{-1}\circ\varphi_{e}$ being a circle homeomorphism called conformal welding. As we have recalled above, this homeomorphism is a quasisymmetry: we deduce from this that $V_{h}:g\mapsto g\circ h$ is an isomorphism of $\mathcal{C}_{p}(\mathbb{T})$ that extends to an isomorphism of $\mathcal{B}_{p,p}^{1/p}(\mathbb{T})$ with $V_{h}^{-1}=V_{h^{-1}}$. Now, if one considers any quasisymmetry $h$ of the unit circle, we know [1] that $h$ is the conformal welding of some quasicircle. We deduce from this geometric approach that $V_{h}$ is an isomorphism of $\mathcal{B}_{p,p}^{1/p}(\mathbb{T})$ for any quasisymmetry $h$. Conversely, Nag-Sullivan [37] for the case $p=2$ and Bourdaud [5] for general case have proven that if $h$ is a circle homeomorphism and if $V_{h}$ is an isomorphism of $\mathcal{B}_{p,p}^{1/p}(\mathbb{T})$ then $h$ must be a quasisymmetry. Vodop’Yanov[48], and later Bourdaud [5] moreover proved that if $s\neq 1/p$ then the circle homeomorphisms $h$ such that $V_{h}$ operates on $\mathcal{B}_{p,p}^{s}(\mathbb{T})$ isomorphically are the bi-Lipschitz ones which form a proper subgroup of the group of quasisymmetries.
2.2 Boundary values of functions in $\mathcal{B}_{p,p}^{1/p}(\Omega)$